Main page‎ > ‎Observing Runs‎ > ‎Night Reports‎ > ‎

Night Report 2016-02-25

Observing: Stefan Taubenberger, Tamar Faran, Markus Kromer

Support Team: N. Elias-Rosa

TAT:  Joe Lyman, Kate Maguire


Weather conditions


Start of night: clear, humidity ~ 30%, wind~ 5 m/s

Middle of night: clear, humidity ~ 30%, wind~ 5 m/s

End of night: clear, humidity ~20%, wind ~ 1 m/s


------------------------------------------------


Calibrations


Evening:

Morning:

------------------------------------------------



00:15 UT: image analysis

00:25 UT: Follow-up: ASASSN-15og, JHK, 400s + 720s + 720s, seeing ~0.7"

00:50 UT: Failure of SOFI *


01:20 UT: Switched to EFOSC

01:30 UT: image analysis

01:40 UT: Classification: Gaia16afr, gr13, 2000s, seeing ~0.9"

02:25 UT: Classification: Gaia16afq, gr13, 2000s, seeing ~0.7"

03:00 UT: Classification: Gaia16afo, gr13, 2000s, seeing ~1.1"

03:45 UT: Classification: Gaia16afe, gr13, 2000s, seeing ~0.8"

04:30 UT: Revised Classification: Gaia16aec, gr13, 3600s, seeing ~0.8"

05:40 UT: Follow-up: PS15ae, Ugriz, 4x100s, 6x200s, 4x100s, 4x100s, 4x100s, seeing ~1"

06:45 UT: Classification: AT2016aqa, gr13, 1500s, seeing ~1"

07:15 UT: Classification: AT2016aqb, gr13, seeing ~0.9"

08:00 UT: Classification: Gaia16afw, gr13, 2000s, seeing ~1"

08:40 UT: Classification: AT2016ahb, gr13, 2100s, seeing ~ 1.1"

09:25 UT: Standard: EG274, gr13, 60s, seeing ~ 1.4"


09:30 UT: End of the night


* The vacuum and cooling of SOFI failed and therefore we had to stop the observations and switch to EFOSC

Comments